OpenStudy (anonymous):

calculus question! please help. integrate: 2sin(0.03t)+1.5

OpenStudy (anonymous):

is that (2*sin(.03))/(1.5) ?????

OpenStudy (anonymous):

no.. it is 2*sin(0.03t) + 1.5

OpenStudy (anonymous):

@satellite73 do you know how to do this? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2/.03cos(.03t)+1.5t+C

OpenStudy (anonymous):

@sbc17 has it

OpenStudy (anonymous):

@sbc17 could you please explain how you got that?

OpenStudy (anonymous):

the anti derivative of \(1.5\) is \(1.5t\) that is clear right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and the 2 is unimportant, it comes right out front of the integral

OpenStudy (anonymous):

so the real question is, what is \[\int\sin(.03t)dt\]

OpenStudy (anonymous):

you know the antiderivative of \(\sin(t)\) is \(-\cos(t)\) so that part is okay yes?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

therefore the real real question is "how do you cope with the \(.03t\) inside the sine and you do that via a mental u - sub in your head you say \(u=.03t\) and so \(du=.03dt\) thus \(\frac{1}{.03}du = dt\) and you intergrate \[\frac{1}{.03}\int\sin(u)du=-\frac{1}{3}\cos(u)=-\frac{1}{.03}\cos(.03t)\]

OpenStudy (anonymous):

an easier example would be \[\int\sin(5t)dt=-\frac{1}{5}\cos(5t)\]

OpenStudy (anonymous):

thank you!